# Large-time behaviour for coupled systems of Lotka-Volterra-type Fokker-Planck equations

### v1

@@ 1 Introduction
In this paper, we study convergence to equilibrium of the solution of a system of Fokker–Planck type equations which describes the evolution of the probability densities ${\mathbf{f}}(x,t)=(f_{1}(x,t),f_{2}(x,t))$, $x\in\mathbb{R}_{+}$, of an interacting particle system characterized by two groups of agents/particles. In the following, this system will feature predator-prey interactions and such that their mean values ${\mathbf{m}}(t)=(m_{1}(t),m_{2}(t))$, given by
@@
namely the classical Lotka-Volterra system with logistic growth [39]. At the macroscopic level, these interactions shape the mutual evolution of this coupled system of particles. A rich literature concentrating on species interacting through space-dependent factors have been developed in recent years, see e.g. [3, 14, 16, 18, 19, 20, 21, 25, 30, 49]. In the direction of interacting systems with cross-diffusion we mention the works [2, 13, 15, 23, 24] and [32, 31, 40] for applications to structured populations. A kinetic approach to such-dynamics have been proposed in [12, 47] where a multi-scale perspective connecting kinetic theory with classical models of population dynamics have been proposed. In particular, in these works a system of mean-field equations have been derived from a microscopic dynamics through the approaches defined in [1, 26, 37]. The system of Fokker–Planck type equations derived in [12] reads
@@
$\sigma_{k}(t)=\sigma_{k}[{\mathbf{m}}(t)],\quad\lambda_{i}(t)=\lambda_{k}[{\mathbf{m}}(t)],\quad\mu_{k}(t)=\mu_{k}[{\mathbf{m}}(t)];\quad k=1,2,$ (4)
characteristics which renders the study of the large–time behaviour of the solutions a challenging problem since the evolution of the mean values is non-monotone in time, in contrast to classical approaches in kinetic theory [4, 17]. System (3) is supplemented with no-flux boundary conditions that guarantee the conservation of total mass.
Relaxation problems for mean-field equations with time-dependent coefficient have been also considered in [5, 6]. Nevertheless, the analysis discussed in [5] has been made possible thanks to the explicit parametrisation of the quasi-equilibrium density which enabled the application of classical entropy methods, see e.g. [27, 28] for related approaches. We recall that, given a Fokker–Planck equation of type (3), namely
@@
The choice of the logistic case is motivated by the fact that, when $\alpha(K-\delta)>0$, the solution $\mathbf{m}(t)$ converges to the unique equilibrium point
@@
In a recent work [35], the large-time behavior of a system of Fokker–Planck type equations with time-dependent diffusion coefficients has been investigated by means of Energy distances, a well-known class of measures of discrepancy between probability distributions, widely used in Statistics [45] and Artificial Intelligence [9]. This analysis suggests that such (weak) distances are particularly well suited for studying the convergence properties of kinetic systems with time-varying parameters. Energy distances, introduced by Székely (cf. [43, 44, 45]), can be regarded as a natural generalization of Cramér’s distance [22], also known as the continuous ranked probability score. They represent a valid alternative to other well-known metrics, such as the Wasserstein distance [48] and the Fourier-based metrics introduced in [29], which have been extensively employed in kinetic theory (see [7] for a recent overview of their properties and mutual relationships). A rigorous validation of the suitability of Energy distances for studying convergence to equilibrium in linear kinetic equations was presented in [8]. There, it was shown that these distances provide an effective tool to quantify the rate of convergence to equilibrium for solutions of linear diffusion equations and Fokker–Planck type equations with linear drift, including the one-dimensional socio-economic models analyzed in [27]. In all considered cases, the solutions exhibit exponential decay in time towards equilibrium.
@@ 3 Lotka-Volterra-type kinetic equations
In detail, let $f_{1}(x,t),f_{2}(x,t)$, $(x,t)\in\mathbb{R}_{+}^{2}$, denote the probability density functions of two types of agents/particles, typically referred to as preys and predators, respectively. Accordingly, $f_{i}(x,t)dx,\,\,i=1,2$ quantifies the fraction of group $i=1,2$ with size in $[x,x+dx]$ at time $t\geq 0$. Moreover, let $m_{1}(t),m_{2}(t)$, as given by (1), the mean values of the populations.
@@ 3.1 Lotka-Volterra-type Fokker-Planck equations
which is meaningful under the hypothesis that $\alpha(K-\delta)>0$. At variance with a purely microscopic approach, the new system of kinetic equations is able to propagate information on the whole moment system of the kinetic distributions $f_{1},f_{2}$. Nevertheless, as investigated in [12], this systems become rapidly unfeasible to direct inspections since nonlinear terms appears starting from the evolution of the second order moment, i.e. $\kappa=2$ in (27). For those reasons, a reduced complexity system of equations have been proposed in [12, 47] building on quasi-invariant limit methods for collisional kinetic equations, see [37] for a comprehensive introduction.
@@
In (LABEL:eq:values) the positive constants $\sigma_{1}^{2},\sigma_{2}^{2}>0$ measure the strength of the diffusion, while the $\theta,\chi>-1$ and $\nu>0$ are related to the rate of birth and death of particles. For large times we have
@@ 3.2 Evolution of observable quantities for the mean-field model
$\begin{split}\dfrac{d}{dt}V_{1}(t)&=\sigma_{2}^{2}(m_{1}+m_{2})m_{1}^{2p}-2\left(\beta m_{2}+\dfrac{\alpha}{K}m_{1}+\alpha\chi\right)V_{1},\\ \dfrac{d}{dt}V_{2}(t)&=\sigma_{2}^{2}m_{1}m^{2p}_{2}-2(\gamma(\mu-m_{1})+\nu\theta)V_{2},\end{split}$
@@
$\begin{split}\dfrac{d}{dt}V_{1}(t)&=-2\left(\beta m_{2}(t)+\dfrac{\alpha}{K}m_{1}(t)+\sigma_{1}^{2}(m_{1}(t)+m_{2}(t))m_{1}(t)\alpha\chi\right)V_{1}(t),\\ \dfrac{d}{dt}V_{2}(t)&=-2(\gamma(\mu-m_{1}(t))+\nu\theta)V_{2}(t)+\sigma_{2}^{2}m_{1}(t)m_{2}(t),\end{split}$ (28)
@@
On the other hand, if $p=1$ we obtain
@@
$2(\beta m_{2}^{\infty}+\frac{\alpha}{K}m_{1}^{\infty}+\alpha\chi)-\sigma_{1}^{2}(m_{1}^{\infty}+m_{2}^{\infty})>0,\qquad 2(\gamma(\mu-m_{1}^{\infty})+\nu\theta)-\sigma_{2}^{2}m_{1}^{\infty}>0$
These explicit results allow to compare the resulting stationary variances, and verify the impact of the parameter $p$. In Figure 1 we represent the evolution of mean and variance from (22)-(28)-(30) for $p=1/2$ and $p=1$. As expected, we may observe how the dynamics of the macroscopic quantities lead to the emergence of a stable configuration of expected values and variances.
@@ 4.1 Quasi-equilibrium distributions
$\begin{split}\bar{\omega}_{1}^{\infty}&=\frac{2\left(\beta m_{2}^{\infty}+\frac{\alpha}{K}m_{1}^{\infty}+\alpha\chi\right)}{\sigma_{1}^{2}(m_{1}^{\infty}+m_{2}^{\infty})},\qquad\bar{\nu}_{1}^{\infty}=\dfrac{2\alpha(\chi+1)m_{1}^{\infty}}{\sigma_{1}^{2}(m_{1}^{\infty}+m_{2}^{\infty})}+1\\ \bar{\omega}_{2}^{\infty}&=\dfrac{2(\gamma(\mu-m_{1}^{\infty})+\nu\theta)}{\sigma_{2}^{2}m_{1}^{\infty}},\quad\qquad\bar{\nu}_{2}^{\infty}=\dfrac{2\nu(\theta+1)m_{2}^{\infty}}{\sigma_{2}^{2}m_{1}^{\infty}}+1,\end{split}$
@@ 4.2 Convergence to equilibrium in Cramér distance
where $M(t)\to 0$ as $t\to+\infty$.
@@ 4.3 Convergence in Energy distance for $p=1/2$ and $p=1$
In this Section, we will prove that in the relevant cases $p=1/2$ and $p=1$ the solutions $f_{k}(x,t)$, $k=1,2$, to the Fokker–Planck equations (40), converge to their equilibria $f_{k}^{\infty}(x,t)$, $k=1,2$ in Energy distance of order $\ell$, provided the parameter $\ell$ characterising the normalised Energy distance (16) is sufficiently big (typically $\ell\geq 1$).
Indeed, the special cases $p=1/2$ and $p=1$ can be treated by writing the Fokker–Planck equations in Fourier transform, which allows a simpler treatment. If the coefficients do not depend on time, the Fokker–Planck equations in (40) belong to the class
$\dfrac{\partial}{\partial t}f(x,t)=\dfrac{\sigma^{2}}{2}\dfrac{\partial^{2}}{\partial x^{2}}(x^{2p}f(x,t))+\dfrac{\partial}{\partial x}((\lambda x-\mu)f(x,t)),$ (50)
where $\sigma^{2},\lambda,\mu$ are time-independent positive parameters. This linear equation is coupled with no-flux boundary conditions in $x=0$. When $p=1/2$, the steady state is the Gamma distribution
@@
$\widehat{h*\kappa_{r}}(x)=\frac{1}{|\xi|^{r}}\hat{h}(\xi).$ (56)
@@
$f^{\infty}(x)=\dfrac{\omega^{\nu}}{\Gamma(\nu)}x^{-1-\nu}\exp\left[\frac{\omega}{x}\right],\qquad\nu=1+\dfrac{2\lambda}{\sigma^{2}},\omega=\frac{2\mu}{\sigma^{2}}.$ (62)
@@
It is immediate to verify that the conclusions of Lemma 2 remain valid even if the coefficients of the Fokker–Planck equation are time dependent, and that in this range of $\ell$ the contribution of the diffusion term is non positive. Hence, we obtain the inequality
$\frac{d}{dt}\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|^{2}}{|\xi|^{2\ell}}\,d\xi$ $\leq-\lambda(t)(2\ell-1)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|^{2}}{|\xi|^{2\ell}}\,d\xi+$ (64)
$2A(t)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|}{|\xi|^{2\ell}}\,\left|\xi\frac{\partial\hat{f}_{k}^{\infty}(\xi)}{\partial\xi}\right|\,d\xi+$
$2B(t)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|}{|\xi|^{2\ell}}\,\left|\xi\hat{f}_{k}^{\infty}(\xi)\right|\,d\xi.$
@@
$\int_{\mathbb{R}}\left|\frac{\partial\hat{f}_{i}^{\infty}(\xi)}{\partial\xi}\right|^{2}\,d\xi=2\pi\int_{\mathbb{R}_{+}}x^{2}|f_{k}^{\infty}(x)|^{2}\,dx,$

### v2

@@ 1 Introduction
In this paper, we study convergence to equilibrium of the solution of a system of Fokker–Planck type equations which describes the evolution of the probability densities ${\mathbf{f}}(x,t)=(f_{1}(x,t),f_{2}(x,t))$, $x\in\mathbb{R}_{+}$, of an interacting particle system characterized by two groups of agents/particles. In the present setting $x\in\mathbb{R}_{+}$ denotes the population density, whose evolution is interpreted in a statistical sense [12, 36]. In the following, this system will feature predator-prey interactions and such that their mean values ${\mathbf{m}}(t)=(m_{1}(t),m_{2}(t))$, given by
@@
where $\alpha>0$ is the growth rate, $K>0$ the carrying capacity, $\beta>0$ the removal rate due to contact between agents in the first population interacting with agents in the second population, $\delta>0$ a natural decay rate in the second specie, and $\gamma>0$ the growth rate of the second population due to interactions between populations. The system (2) corresponds to the classical Lotka-Volterra model with logistic growth [40]. At the macroscopic level, these interactions shape the mutual evolution of this coupled system of particles. A rich literature concentrating on species interacting through space-dependent factors have been developed in recent years, see e.g. [3, 14, 16, 18, 19, 20, 21, 25, 30, 50]. In the direction of interacting systems with cross-diffusion we mention the works [2, 13, 15, 23, 24] and [31, 32, 41] for applications to structured populations. A kinetic approach to such-dynamics have been proposed in [12, 48] where a multi-scale perspective connecting kinetic theory with classical models of population dynamics have been proposed. In particular, in these works a system of mean-field equations have been derived from a microscopic dynamics through the approaches defined in [1, 26, 38]. The system of Fokker–Planck type equations derived in [12] reads
@@
$\sigma_{k}(t)=\sigma_{k}[{\mathbf{m}}(t)],\quad{\lambda_{k}}(t)=\lambda_{k}[{\mathbf{m}}(t)],\quad\mu_{k}(t)=\mu_{k}[{\mathbf{m}}(t)];\quad k=1,2,$ (4)
characteristics which renders the study of the large–time behaviour of the solutions a challenging problem since the evolution of the mean values is non-monotone in time, in contrast to classical approaches in kinetic theory [4, 17]. We recall that the bound $1/2\leq p\leq 1$ is linked to the integrability of equilibrium densities. System (3) is supplemented with no-flux boundary conditions that guarantee the conservation of total mass.
Relaxation problems for mean-field equations with time-dependent coefficient have been also considered in [5, 6]. Nevertheless, the analysis discussed in [5] has been made possible thanks to the explicit parametrisation of the quasi-equilibrium density which enabled the application of classical entropy methods, see e.g. [27, 28, 33] for related approaches. We recall that, given a Fokker–Planck equation of type (3), namely
@@
The choice of the logistic case is motivated by the fact that, when $(\gamma K-\delta)>0$}, the solution $\mathbf{m}(t)$ converges to the unique equilibrium point
@@
In a recent works [35, 36], the large-time behaviour of a system of Fokker–Planck type equations with time-dependent diffusion coefficients has been investigated by means of Energy distances, a well-known class of measures of discrepancy between probability distributions, widely used in Statistics [46] and Artificial Intelligence [9]. This analysis suggests that such (weak) distances are particularly well suited for studying the convergence properties of kinetic systems with time-varying parameters. Energy distances, introduced by Székely (cf. [44, 45, 46]), can be regarded as a natural generalization of Cramér’s distance [22], also known as the continuous ranked probability score. They represent a valid alternative to other well-known metrics, such as the Wasserstein distance [49] and the Fourier-based metrics introduced in [29], which have been extensively employed in kinetic theory (see [7] for a recent overview of their properties and mutual relationships). A rigorous validation of the suitability of Energy distances for studying convergence to equilibrium in linear kinetic equations was presented in [8]. There, it was shown that these distances provide an effective tool to quantify the rate of convergence to equilibrium for solutions of linear diffusion equations and Fokker–Planck type equations with linear drift, including the one-dimensional socio-economic models analysed in [27]. In all considered cases, the solutions exhibit exponential decay in time towards equilibrium.
@@ 3 Lotka-Volterra-type kinetic equations
In detail, let $f_{1}(x,t),f_{2}(x,t)$, $(x,t)\in\mathbb{R}_{+}^{2}$, denote the probability density functions of two types of agents/particles, typically referred to as preys and predators, respectively. Accordingly, $f_{k}(x,t)dx,\,\,k=1,2$ quantifies the fraction of group $k=1,2$ with size in $[x,x+dx]$ at time $t\geq 0$. Moreover, let $m_{1}(t),m_{2}(t)$, as given by (1), the mean values of the populations.
@@ 3.1 Lotka-Volterra-type Fokker-Planck equations
which is meaningful under the hypothesis that $\gamma K-\delta>0$. At variance with a purely microscopic approach, the new system of kinetic equations is able to propagate information on the whole moment system of the kinetic distributions $f_{1},f_{2}$ provided by the quantities
$m_{k}^{(\ell)}=\int_{\mathbb{R}_{+}}x^{\ell}f_{k}(x,t)dx,\qquad\ell>0,k=1,2$ (24)
Nevertheless, as investigated in [12], this systems become rapidly unfeasible to direct inspections since nonlinear terms appears starting from the evolution of the second order moment, i.e. $\ell=2$ in (24). For those reasons, a reduced complexity system of equations have been proposed in [12, 48] building on quasi-invariant limit methods for collisional kinetic equations, see [38] for a comprehensive introduction.
@@
In (LABEL:eq:values) the positive constants $\sigma_{1}^{2},\sigma_{2}^{2}>0$ measure the strength of the diffusion, while the $\theta,\chi>-1$ and $\nu>0$ are related to the rate of birth and death of particles. Furthermore, according to [12] the parameter $\delta>0$ is such that $\delta=\gamma\mu-\nu$. For large times we have
@@ 3.2 Evolution of observable quantities for the mean-field model
$\begin{split}\dfrac{d}{dt}V_{1}(t)&={\sigma_{1}^{2}}(m_{1}+m_{2})m_{1}^{{(2p)}}-2\left(\beta m_{2}+\dfrac{\alpha}{K}m_{1}+\alpha\chi\right)V_{1},\\ \dfrac{d}{dt}V_{2}(t)&=\sigma_{2}^{2}m_{1}m^{{(2p)}}_{2}-2(\gamma(\mu-m_{1})+\nu\theta)V_{2},\end{split}$
@@
$\begin{split}\dfrac{d}{dt}V_{1}(t)&={-2\left(\beta m_{2}(t)+\dfrac{\alpha}{K}m_{1}(t)+\alpha\chi\right)V_{1}(t)+\sigma_{1}^{2}(m_{1}(t)+m_{2}(t))m_{1}(t)}\\ \dfrac{d}{dt}V_{2}(t)&=-2(\gamma(\mu-m_{1}(t))+\nu\theta)V_{2}(t)+\sigma_{2}^{2}m_{1}(t)m_{2}(t),\end{split}$ (29)
@@
where $m_{1}^{\infty},m_{2}^{\infty}$ have been defined in (23). On the other hand, if $p=1$ we obtain
@@
$2(\beta m_{2}^{\infty}+\frac{\alpha}{K}m_{1}^{\infty}+\alpha\chi)-\sigma_{1}^{2}(m_{1}^{\infty}+m_{2}^{\infty})>0,\qquad 2(\gamma(\mu-m_{1}^{\infty})+\nu\theta)-\sigma_{2}^{2}m_{1}^{\infty}>0,$
which, taking into account the equilibrium values (23), imply a bound from above on the diffusion coefficients given by
$\sigma_{1}^{2}<\dfrac{\delta\beta K+\alpha(\gamma K-\delta)}{2\alpha\beta\gamma K(\chi+1)},\qquad\sigma_{2}^{2}<\dfrac{2\gamma(\gamma\mu-\delta+\nu\theta)}{\delta}.$
We may observe that, under the hypotheses discussed in [12], the upper bound for $\sigma_{1}^{2}$ is always positive while the one connected to $\sigma_{2}^{2}$ is positive under the identification $\delta=\gamma\mu-\nu$. These explicit results allow to compare the resulting stationary variances, and verify the impact of the parameter $p$. In Figure 1 we represent the evolution of mean and variance from (22)-(29)-(31) for $p=1/2$ and $p=1$. As expected, we may observe how the dynamics of the macroscopic quantities lead to the emergence of a stable configuration of expected values and variances.
@@ 4.1 Quasi-equilibrium distributions
$\begin{split}\bar{\omega}_{1}^{\infty}&=\frac{2\left(\beta m_{2}^{\infty}+\frac{\alpha}{K}m_{1}^{\infty}+\alpha\chi\right)}{\sigma_{1}^{2}(m_{1}^{\infty}+m_{2}^{\infty})},\qquad\bar{\nu}_{1}^{\infty}=\dfrac{2\alpha(\chi+1)m_{1}^{\infty}}{\sigma_{1}^{2}(m_{1}^{\infty}+m_{2}^{\infty})}+1\\ \bar{\omega}_{2}^{\infty}&=\dfrac{2(\gamma(\mu-m_{1}^{\infty})+\nu\theta)}{\sigma_{2}^{2}m_{1}^{\infty}},\quad\qquad\bar{\nu}_{2}^{\infty}=\dfrac{2\nu(\theta+1)m_{2}^{\infty}}{\sigma_{2}^{2}m_{1}^{\infty}}+1,\end{split}$ (41)
@@ 4.2 Convergence to equilibrium in Cramér distance
where $M(t)\to 0$ as $t\to+\infty$ and $\lambda_{k}(t)\geq\lambda^{*}>0$, $k=1,2$, for all $t\geq 0$.
@@
Remark 3.
From (23) we conclude that $\lambda_{1,\infty},\lambda_{2,\infty}>0$. Since $m_{1}(t)\to m_{1}^{\infty}$, $m_{2}(t)\to m_{2}^{\infty}$ as $t\to+\infty$, there exists $t^{*}>0$ such that the condition $\lambda_{1}(t),\lambda_{2}(t)\geq\lambda^{*}>0$ is true for any $t>t^{*}$.
@@ 4.3 Convergence in Energy distance for $p=1/2$ and $p=1$
In this Section, we will prove that in the relevant cases $p=1/2$ and $p=1$ the solutions $f_{k}(x,t)$, $k=1,2$, to the Fokker–Planck equations (42), converge to their equilibria $f_{k}^{\infty}(x)$, $k=1,2$ in Energy distance of order $\ell$, provided the parameter $\ell$ characterising the normalised Energy distance (16) is sufficiently big (typically $\ell\geq 1$). Indeed, the special cases $p=1/2$ and $p=1$ can be treated by writing the Fokker–Planck equations in Fourier transform, which allows direct computations. In the following we first study the case in which coefficients do not depend on time.
4.3.1 The time-independent case
If the coefficients do not depend on time, i.e. $\sigma^{2},\lambda,\mu>0$ are time-independent positive parameter, the Fokker–Planck equations in (42) belong to the class
$\dfrac{\partial}{\partial t}f(x,t)=\dfrac{\sigma^{2}}{2}\dfrac{\partial^{2}}{\partial x^{2}}(x^{2p}f(x,t))+\dfrac{\partial}{\partial x}((\lambda x-\mu)f(x,t)).$ (52)
This linear equation is coupled with no-flux boundary conditions in $x=0$. When $p=1/2$, the steady state is the Gamma distribution
@@ 4.3.1 The time-independent case
$\widehat{h*\kappa_{r}}{(\xi)}=\frac{1}{|\xi|^{r}}\hat{h}(\xi).$ (58)
@@
$f^{\infty}(x)=\dfrac{\omega^{\nu}}{\Gamma(\nu)}x^{-1-\nu}\exp{\left[-\frac{\omega}{x}\right]},\qquad\nu=1+\dfrac{2\lambda}{\sigma^{2}},\omega=\frac{2\mu}{\sigma^{2}}.$ (64)
@@ 4.3.2 The time dependent case
It is immediate to verify that the conclusions of Lemma 2 remain valid even if the coefficients of the Fokker–Planck equation are time dependent, and that in this range of $\ell$ the contribution of the diffusion term is non positive. Hence, the results of the time-independent case provided in Section 4.3.1 lead to the following inequality if $p=1/2$
$\frac{d}{dt}\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|^{2}}{|\xi|^{2\ell}}\,d\xi$ $\leq-{\lambda_{k}}(t)(2\ell-1)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|^{2}}{|\xi|^{2\ell}}\,d\xi+$ (66)
$2{A_{k}}(t)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|}{|\xi|^{2\ell}}\,\left|\xi\frac{\partial\hat{f}_{k}^{\infty}(\xi)}{\partial\xi}\right|\,d\xi+$
$2{B_{k}}(t)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|}{|\xi|^{2\ell}}\,\left|\xi\hat{f}_{k}^{\infty}(\xi)\right|\,d\xi.$
@@
$\int_{\mathbb{R}}\left|\frac{\partial\hat{f}_{k}^{\infty}(\xi)}{\partial\xi}\right|^{2}\,d\xi=2\pi\int_{\mathbb{R}_{+}}x^{2}|f_{k}^{\infty}(x)|^{2}\,dx,$
@@
Analogous results are obtained if $p=1$, for which we get
$\frac{d}{dt}\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|^{2}}{|\xi|^{2\ell}}\,d\xi$ (73)
$\quad\leq-\left[\sigma^{2}_{k}(t)\dfrac{3-2\ell}{4}+{\lambda_{k}}(t)\right](2\ell-1)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|^{2}}{|\xi|^{2\ell}}\,d\xi+$
$\qquad 2{A_{k}}(t)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|}{|\xi|^{2\ell}}\,\left|\xi\frac{\partial\hat{f}_{k}^{\infty}(\xi)}{\partial\xi}\right|\,d\xi+$
$\qquad 2{B_{k}}(t)\int_{\mathbb{R}}\frac{|\hat{f}_{k}(\xi)-\hat{f}_{k}^{\infty}(\xi)|}{|\xi|^{2\ell}}\,\left|\xi\hat{f}_{k}^{\infty}(\xi)\right|\,d\xi.$
@@
Remark 4.
We observe that if $p=1/2$ the functions $f_{k}^{\infty}(x)$ and $xf_{k}^{\infty}(x)$ are always in $L^{2}(\mathbb{R}_{+})$ since all the moments are finite. In the case $p=1$ the quantities $f_{k}^{\infty}(x)$ and $xf_{k}^{\infty}(x)$ are in $L^{2}(\mathbb{R}_{+})$ under the hypotesis of bounded variance, holding under the assumptions $\bar{\nu}^{\infty}_{k}>3$, see (41).